Q.9

The balance sheet for Solomon Corporation follows:

Current assets $ 237,000
Long-term assets (net) 752,000
Total assets $ 989,000
Current liabilities $ 156,000
Long-term liabilities 450,000
Total liabilities 606,000
Common stock and retained earnings 383,000
Total liabilities and stockholders’ equity $ 989,000


Required

Compute the following. (Round "Ratios" to 1 decimal place.)

Working capital Current ratio Debt-to-assets ratio Debt-to-equity ratio
Answer #1

Working capital = Current assets - Current Liabilities

= 237,000 - 156000

= 81,000

Current Ratio = Current assets /Current liabilities

= 237000/156000

= 1.5

Debt to assets ratio = Total Liabilities /Total assets

= 606,000/989,000

= 0.6

Debt to equity = Total Liabilities /Total equity

= 606,000/383,000

= 1.6

Comment if you face any issues